OpenStudy (anonymous):

solve. check for extraneous solutions the square root of x + 8 equals x - 4

OpenStudy (mertsj):

Square both sides.

OpenStudy (anonymous):

how do i do that?

OpenStudy (mertsj):

\[(\sqrt{x+8})^2=(x-4)^2\]

OpenStudy (mertsj):

\[(\sqrt{x+8})(\sqrt{x+8})=(x-4)(x-4)\]

OpenStudy (anonymous):

which would be x + 64 = x + 16?

OpenStudy (mertsj):

No. On the left use this property:

OpenStudy (mertsj):

\[(\sqrt{x})^2=x\]

OpenStudy (mertsj):

On the right use this property:

OpenStudy (mertsj):

\[(a-b)(a-b)=a^2-2ab+b^2\]

OpenStudy (anonymous):

x^2 - 2x8 + 64?

OpenStudy (mertsj):

x^2-8x+16

OpenStudy (anonymous):

i don't understand, =/

OpenStudy (mertsj):

\[(\sqrt{x+8})(\sqrt{x+8})=x+8\]

OpenStudy (mertsj):

\[(x-4)(x-4)=x^2-4x-4x+16=x^2-8x+16\]

OpenStudy (mertsj):

So now we have: \[x^2-8x+16=x+8\]

OpenStudy (mertsj):

Subtract x from both sides. Subtract 8 from both sides.

OpenStudy (mertsj):

Then factor.